ADNAN MUHAMMED

Grade 12,

If (1+3+5+.....+p) + (1+3+5+.....+q) = (1+3+5+.....+r) where each set of parentheses contains the sum of consecutive integers as shown, the smallest possible value of p+q+r, (where p>6) is (a) 12 (b) 21 (c) 45 (d) 54

If (1+3+5+.....+p) + (1+3+5+.....+q) = (1+3+5+.....+r) where each set of parentheses contains the sum of consecutive integers as shown, the smallest possible value of p+q+r, (where p>6) is


(a) 12


(b) 21


(c) 45


(d) 54

Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
12 years ago

Dear student,

The correct option is c)

BEST OF LUCK..!!!!


 

We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh - IIT Delhi

iit jee
11 years ago

sagar singh....u r mad....

ans is 21.

i''ll show u how...--->

no. of terms--> (p+1)/2, (q+1)/2, (r+1)/2

sum of 1st n odd nos. =n^2

 ((p+1)/2)^2 +  ((q+1)/2)^2 =  ((r+1)/2)^2

(p+1)^2 + (q+1)^2 = (r+1)^2

p+1,q+1,r+1 are pythagorean triplets

p>6...........p+1>7

therfore the smallest set is (8,6,10)......p+1=8,q+1=6,r+1=10................p+q+r=21......ans....simple